if grand son is of one day then father is of one week and if grandson is of 1 month then grandfather is of 1 year if the total of their ages is 140 years so whats their ages
Answer by ankor@dixie-net.com(22740)

If grand son is of one day, then father is of one week.
And if grandson is of 1 month, then grandfather is of 1 year.
If the total of their ages is 140 years, so what's their ages.
:
From the information given, we can say:
:
The grandson to father ratio: 1:7
The grandson to grandfather ratio: 1:12
:
Let x = the multiplier
:
1x + 7x + 12x = 140
20x = 140
x = 140/20
x = 7 is the multiplier
:
Grandson is 7*1 = 7 yrs old
Father is 7*7 = 49 yrs old
Grandfather is 7*12 = 84 yrs old
:
:
Check 7 + 49 + 84 = 140